OpenStudy (anonymous):

How long will it take for $2000 to double if it is invested at 13% interest compounded continuously?

OpenStudy (anonymous):

Assuming compounded pa, 4000 = 2000(1.13)^t Somebody verify this.

OpenStudy (anonymous):

4000=2000(1.13)^t (2000(1.13)^t)/2000=4000/2000 (1.13)^t=2 2yrs

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

2 is wrong. Its \[\log_{1.13} 2 \]

OpenStudy (anonymous):

that is 5.67 years

OpenStudy (anonymous):

i dont understand how you got log1.13^2

OpenStudy (anonymous):

its not 1.13^2 its base its like \[\log_{1.13} 2\]

OpenStudy (anonymous):

1.13 is the base

OpenStudy (anonymous):